OpenStudy (lukecrayonz):

find the nth partial sum of the arithmetic sequence 1.50, 1.45, 1.40, 1.35, ...., n=20

OpenStudy (anonymous):

Following the pattern, the sequence of partial sums can be expressed as: \[s _{n} = 1.55 - n(0.05)\] since we are looking for n=20, we plug that into our sequence formula: \[s_{20} = 1.55 - (20)(0.05) = 1.55 - 1 = 0.55 \]

OpenStudy (anonymous):

difference= -.05 (1.45-1.50) so the formula for arithmetic series is Tn=t1+(n−1)d . Where T1 is the first term and d is the difference. So, T20= 1.50 + (20-1)(-.05) --> T20 = 1.50 -0.95 Therefore, T20 is 0.55

OpenStudy (asnaseer):

@Lukecrayonz: both @Hermeezey and @dianneg have shown how to get the 20'th term of this sequence. however, the question is asking for the sum of the first 20 terms. so you need to use the formula for the sum of an arithmetic series, which is:\[S_n=\frac{n}{2}(2a_1+(n-1)d)\]where: \(S_n\) = sum of the first n terms n = number of terms \(a_1\) = the first term d = common difference between each term

OpenStudy (anonymous):

haha @asnaseer you're right, I skimmed the question and mis-answered it.

OpenStudy (anonymous):

Yeah, so did I. LOL. my bad.